OpenStudy (anonymous):

At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path.The bicyclist heading north is riding 6 km/hour faster than the bicyclist heading south. At 10:15, they are 42.5 km apart. Find the two bicyclists’ rates.

OpenStudy (anonymous):

\[3\frac{1}{4}x +3\frac{1}{4}(x + 6) = 42.5\] \[2x + 6 = \frac{425}{10} \times \frac{4}{3} \]